# presep= ]e1,e2,e3

### v1

@@ A.3 Proof of Theorem 3 (Hill consistency)
Let $\varepsilon>0$ be given and assume that $q_{n}/u_{n}>1+\varepsilon$ for all $n\in\mathbb{N}$. Since $\overline{F}^{x}_{n}$ is decreasing we have
$\widetilde{T}^{x}_{n}\left(1+\varepsilon\right)=\frac{\overline{F}^{x}_{n}(\left(1+\varepsilon\right)u_{n})}{\overline{F}^{x}(u_{n})}\geq\frac{\overline{F}^{x}_{n}(q_{n})}{\overline{F}^{x}(u_{n})}=\frac{\overline{F}^{x}_{n}(F^{\leftarrow,x}_{n}\left(1-k_{n}/n\right))}{\overline{F}^{x}(F^{\leftarrow,x}\left(1-k_{n}/n\right))}=1.$
@@
$\mathrm{Pr}\left(\frac{q_{n}}{u_{n}}>1+\varepsilon\right)\leq\mathrm{Pr}\left(\widetilde{T}^{x}_{n}\left(1+\varepsilon\right)>1\right).$
Since (4) holds, then $\widetilde{T}^{x}_{n}\left(1+\varepsilon\right)\overset{\mathrm{Pr}}{\to}\left(1+\varepsilon\right)^{-\alpha_{x}}$ and thus
$\limsup_{n\to\infty}\mathrm{Pr}\left(\frac{q_{n}}{u_{n}}>1+\varepsilon\right)\leq\lim_{n\to\infty}\mathrm{Pr}\left(\widetilde{T}^{x}_{n}\left(1+\varepsilon\right)>1\right)=0.$
Using similar arguments we obtain that $\limsup_{n\to\infty}\mathrm{Pr}\left({q_{n}}/{u_{n}}<1-\varepsilon\right)=0$. ∎
@@
$\widehat{\gamma}_{n}\left(x\right)=\int_{q_{n}/u_{n}}^{1}\frac{\widetilde{T}_{n}^{x}\left(s\right)}{s}\mathrm{d}s+\int_{1}^{\infty}\frac{\widetilde{T}_{n}^{x}\left(s\right)}{s}\mathrm{d}s.$
Proposition 1 yields that $q_{n}/u_{n}\overset{\mathrm{Pr}}{\to}1$ and thus the first integral is $o_{\mathrm{Pr}}\left(1\right)$.
The convergence (4) is uniform on compacts of $[1,\infty)$ and hence for all $t>1$ we have that
@@
$=\frac{1}{\eta}\int_{t}^{\infty}\frac{\overline{F}^{x}(su_{n})}{s\overline{F}^{x}(u_{n})}\int(su_{n})^{-(\alpha_{x-uh_{n}}-\alpha_{x})}\frac{L^{x-uh_{n}}(su_{n})}{L^{x}(su_{n})}\frac{g(x-h_{n}u)}{g(x)}K\left(u\right)\mathrm{d}u\mathrm{d}s$
$\leq\frac{1}{\eta}\int_{t}^{\infty}\frac{\overline{F}^{x}(su_{n})}{s\overline{F}^{x}(u_{n})}\sup_{u\in[-1,1]}\left((su_{n})^{-(\alpha_{x-uh_{n}}-\alpha_{x})}\frac{L^{x-uh_{n}}(su_{n})}{L^{x}(su_{n})}\frac{g(x-h_{n}u)}{g(x)}\right)\mathrm{d}s$
$\leq\frac{1}{\eta}\int_{t}^{\infty}\frac{\overline{F}^{x}(su_{n})}{s\overline{F}^{x}(u_{n})}\sup_{u\in[-1,1]}\left(\frac{g(x-h_{n}u)}{g(x)}\right)s^{\left(c_{\alpha}+c_{L}\right)h_{n}}u_{n}^{\left(c_{\alpha}+c_{L}\right)h_{n}}\mathrm{d}s$
$\leq\frac{1}{\eta}\sup_{u\in[-1,1]}\left(\frac{g(x-h_{n}u)}{g(x)}\right)u_{n}^{\left(c_{\alpha}+c_{L}\right)h_{n}}\int_{t}^{\infty}s^{-\alpha_{x}+\varepsilon-1+\left(c_{\alpha}+c_{L}\right)h_{n}}\mathrm{d}s,$
for $\varepsilon\in(0,\alpha_{x})$, using Potter’s bound. Let $\varepsilon^{\prime}\in(0,\alpha_{x}-\varepsilon)$ be given. It is possible to find $n^{\prime}$ such that $\left(c_{\alpha}+c_{L}\right)h_{n}\leq\varepsilon^{\prime}$ for all $n\geq n^{\prime}$. Hence for $n\geq n^{\prime}$ we have that the integrand $s\mapsto s^{-\alpha_{x}+\varepsilon-1+\left(c_{\alpha}+c_{L}\right)h_{n}}$ is integrable and consequently
@@
$\leq\limsup_{n\to\infty}\left\{\frac{1}{\eta}\sup_{u\in[-1,1]}\left(\frac{g(x-h_{n}u)}{g(x)}\right)u_{n}^{\left(c_{\alpha}+c_{L}\right)h_{n}}\int_{t}^{\infty}s^{-\alpha_{x}+\varepsilon-1+\left(c_{\alpha}+c_{L}\right)h_{n}}\right\}=\mathcal{O}\left(t^{-\alpha_{x}+\varepsilon+\varepsilon^{\prime}}\right),$
@@ A.4 Proof of Theorem 4 (Functional CLT)
are mutually independent and each have the same distribution as the first original sample block $\left\{\left(X_{1},Y_{1}\right),\ldots,\left(X_{r_{n}},Y_{r_{n}}\right)\right\}$.
@@
$\widetilde{V}^{{\dagger},x}_{n}(s)=\frac{1}{nh_{n}\overline{F}^{x}(u_{n})}\sum_{i=1}^{m_{n}}\sum_{j=1}^{r_{n}}K\left(\frac{x-X^{{\dagger}}_{(i-1)r_{n}+j}}{h_{n}}\right)1_{\left\{Y^{{\dagger}}_{(i-1)r_{n}+j}>su_{n}\right\}},$
@@
$\lim_{n\to\infty}\operatorname{cov}{\left(\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(s),\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(t)\right)}=\left(s\vee t\right)^{-\alpha_{x}}g(x)\int K^{2}(u)\mathrm{d}u.$ (16)
@@
$\operatorname{cov}{\left(\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(s),\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(t)\right)}=\frac{m_{n}r_{n}}{nh_{n}\overline{F}^{x}(u_{n})}\biggl{[}\mathbb{E}\left[1_{\left\{Y_{0}>\left(s\vee t\right)u_{n}\right\}}K^{2}\left(\frac{x-X_{0}}{h_{n}}\right)\right]$
$\quad\quad-r_{n}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}\right]\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}\right\}}\right]$
$\quad\quad+\sum_{j=1}^{r_{n}}\left(1-\frac{j}{r_{n}}\right)\biggl{\{}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}\right\}}\right]$
$\qquad\qquad\qquad\qquad\qquad+\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right]\biggr{\}}\biggr{]}.$
@@
$\lim_{n\to\infty}\frac{m_{n}r_{n}}{n}\frac{\mathbb{E}\left[1_{\left\{Y_{0}>\left(s\vee t\right)u_{n}\right\}}K^{2}\left(\frac{x-X_{0}}{h_{n}}\right)\right]}{h_{n}\overline{F}^{x}(u_{n})}=\left(s\vee t\right)^{-\alpha_{x}}g(x)\int K^{2}\left(u\right)\mathrm{d}u,$
which is the asymptotic covariance. It remains to show that the second and third terms disappear asymptotically. The second term vanishes by the assumption $\lim_{n\to\infty}{r_{n}h_{n}\overline{F}^{x}(u_{n})}=0$. Since $m_{n}r_{n}/n=\mathcal{O}(1),$, the leading term in the third term is asymptotically proportional to
$\frac{1}{h_{n}\overline{F}^{x}(u_{n})}\sum_{j=1}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right].$
@@
$\limsup_{n\to\infty}\frac{1}{h_{n}\overline{F}^{x}(u_{n})}\sum_{j=1}^{m-1}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right]$
$=\limsup_{n\to\infty}\sum_{j=1}^{m-1}h_{n}\int_{[-1,1]^{2}}\biggl{\{}\frac{\mathbb{E}\left[1_{\left\{Y_{0}>tu_{n}\right\}}1_{\left\{Y_{j}>su_{n}\right\}}\mid X_{0}=x+w_{1}h_{n},X_{j}=x+w_{2}h_{n}\right]}{\overline{F}^{x}(u_{n})}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times g(x+w_{1}h_{n},x+w_{2}h_{n})K\left(w_{1}\right)K\left(w_{2}\right)\biggr{\}}\mathrm{d}(w_{1},w_{2})=0.$
@@
$\limsup_{n\to\infty}\frac{1}{h_{n}\overline{F}^{x}(u_{n})}\sum_{j=m}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right]\leq\varepsilon.$
@@
Assume conditions 1-4, 1-2. Then the finite dimensional distributions of $\widetilde{\mathbb{V}}^{{\dagger},x}_{n}$ converge to those of a Gaussian process $\mathbb{V}^{x}$ with covariance function defined in (16).
@@
$S_{n,i}^{x}(s)=\sum_{j=1}^{r_{n}}K\left(\frac{x-X_{(i-1)r_{n}+j}}{h_{n}}\right)1_{\left\{Y_{(i-1)r_{n}+j}>su_{n}\right\}},\quad\overline{S}_{n,i}^{x}(s)=S_{n,i}^{x}(s)-\mathbb{E}\left[S_{n,1}^{x}(s)\right],$
$S_{n,i}^{{\dagger},x}(s)=\sum_{j=1}^{r_{n}}K\left(\frac{x-X^{{\dagger}}_{(i-1)r_{n}+j}}{h_{n}}\right)1_{\left\{Y^{{\dagger}}_{(i-1)r_{n}+j}>su_{n}\right\}},\quad\overline{S}_{n,i}^{{\dagger},x}(s)=S_{n,i}^{{\dagger},x}(s)-\mathbb{E}\left[S_{n,1}^{x}(s)\right],$
@@
where the summands are i.i.d. In the calculations below we show that the conditions of the Lindeberg–Lévy CLT are fulfilled. By Lemma 6 it only remains to show the negligibility condition. Since $\overline{S}_{n,1}^{x}(s)$ and $\overline{S}_{n,1}^{{\dagger},x}(s)$ have the same distribution, we may establish it for the former. Note that $m_{n}/(nh_{n}\overline{F}^{x}(u_{n}))=\mathcal{O}((r_{n}h_{n}\overline{F}^{x}(u_{n}))^{-1})$. Thus the negligibility condition thus follows if the expectation below is $o\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right)$:
@@
$\mathbb{E}\left[\left(\sum_{j=1}^{r_{n}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}\right)^{2}1_{\left\{\lvert\overline{S}_{n,1}^{x}(s)\rvert>\varepsilon\sqrt{k_{n}h_{n}}\right\}}\right]$
$\quad+\left(r_{n}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}\right]\right)^{2}\mathrm{Pr}{\left(\left\lvert\overline{S}_{n,1}^{x}(s)\right\rvert>\varepsilon\sqrt{k_{n}h_{n}}\right)}$
$\quad-2r_{n}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}\right]\sum_{j=1}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}1_{\left\{\lvert\overline{S}_{n,1}^{x}(s)\rvert>\varepsilon\sqrt{k_{n}h_{n}}\right\}}\right].$
@@
$r_{n}^{2}\ \mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}\right]^{2}=r_{n}^{2}\mathcal{O}\left(\left(h_{n}\overline{F}^{x}(u_{n})\right)^{2}\right)=o\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right).$
@@
$I_{1}$ $\leq\frac{1}{\varepsilon\sqrt{k_{n}h_{n}}}\sum_{j=1}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)^{2}1_{\left\{Y_{j}>su_{n}\right\}}\left\lvert\overline{S}_{n,1}^{x}(s)\right\rvert\right]$
$\leq\frac{1}{\varepsilon\sqrt{k_{n}h_{n}}}\sum_{j=1}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)^{2}1_{\left\{Y_{j}>su_{n}\right\}}\left(S_{n,1}^{x}(s)+\mathbb{E}\left[S_{n,1}^{x}(s)\right]\right)\right]$
@@
$\frac{\mathcal{O}\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right)}{\varepsilon\sqrt{k_{n}h_{n}}}\sum_{j=1}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)^{2}1_{\left\{Y_{j}>su_{n}\right\}}\right]=\frac{\mathcal{O}\left(\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right)^{2}\right)}{\varepsilon\sqrt{k_{n}h_{n}}}=o\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right),$

### v2

@@ A.3 Proof of Theorem 3 (Hill consistency)
Let $\varepsilon>0$ be given and assume that $q_{n,k_{n}}^{x}/u_{n}^{x}>1+\varepsilon$ for all $n\in\mathbb{N}$. Since $\overline{F}^{x}_{n}$ is decreasing we have
$\widetilde{T}^{x}_{n}\left(1+\varepsilon\right)=\frac{\overline{F}^{x}_{n}(\left(1+\varepsilon\right)u_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}\geq\frac{\overline{F}^{x}_{n}(q_{n,k_{n}}^{x})}{\overline{F}^{x}(u_{n}^{x})}=\frac{\overline{F}^{x}_{n}(F^{\leftarrow,x}_{n}\left(1-k_{n}/n\right))}{\overline{F}^{x}(F^{\leftarrow,x}\left(1-k_{n}/n\right))}=1.$
@@
$\mathrm{Pr}\left(\frac{q_{n,k_{n}}^{x}}{u_{n}^{x}}>1+\varepsilon\right)\leq\mathrm{Pr}\left(\widetilde{T}^{x}_{n}\left(1+\varepsilon\right)>1\right).$
Since (2) holds, then $\widetilde{T}^{x}_{n}\left(1+\varepsilon\right)\overset{\mathrm{Pr}}{\to}\left(1+\varepsilon\right)^{-1/\gamma(x)}$ and thus
$\limsup_{n\to\infty}\mathrm{Pr}\left(\frac{q_{n,k_{n}}^{x}}{u_{n}^{x}}>1+\varepsilon\right)\leq\lim_{n\to\infty}\mathrm{Pr}\left(\widetilde{T}^{x}_{n}\left(1+\varepsilon\right)>1\right)=0.$
Using similar arguments we obtain that $\limsup_{n\to\infty}\mathrm{Pr}({q_{n,k_{n}}^{x}}/{u_{n}^{x}}<1-\varepsilon)=0$. ∎
@@
$\widehat{\gamma}_{n}\left(x\right)=\int_{q_{n,k_{n}}^{x}/u_{n}^{x}}^{1}\frac{\widetilde{T}_{n}^{x}\left(s\right)}{s}\mathrm{d}s+\int_{1}^{\infty}\frac{\widetilde{T}_{n}^{x}\left(s\right)}{s}\mathrm{d}s.$
Proposition 1 yields that $q_{n,k_{n}}^{x}/u_{n}^{x}\overset{\mathrm{Pr}}{\to}1$ and thus the first integral is $o_{\mathrm{Pr}}\left(1\right)$.
The convergence of equation (2) is uniform on compacts of $1,\infty)$ and hence for all $t>1$ we have that
@@
$=\frac{1}{\eta}\int_{t}^{\infty}\frac{\overline{F}^{x}(su_{n}^{x})}{s\overline{F}^{x}(u_{n}^{x})}\int(su_{n}^{x})^{-(1/\gamma(x-uh_{n})-1/\gamma(x))}\frac{L^{x-uh_{n}}(su_{n}^{x})}{L^{x}(su_{n}^{x})}\frac{g(x-h_{n}u)}{g(x)}K\left(u\right)\mathrm{d}u\mathrm{d}s$
$\leq\frac{1}{\eta}\int_{t}^{\infty}\frac{\overline{F}^{x}(su_{n}^{x})}{s\overline{F}^{x}(u_{n}^{x})}\sup_{u\in[-1,1]}\left((su_{n}^{x})^{-(1/\gamma(x-uh_{n})-1/\gamma(x))}\frac{L^{x-uh_{n}}(su_{n}^{x})}{L^{x}(su_{n}^{x})}\frac{g(x-h_{n}u)}{g(x)}\right)\mathrm{d}s$
$\leq\frac{1}{\eta}\int_{t}^{\infty}\frac{\overline{F}^{x}(su_{n}^{x})}{s\overline{F}^{x}(u_{n}^{x})}\sup_{u\in[-1,1]}\left(\frac{g(x-h_{n}u)}{g(x)}\right)s^{\left(c_{\gamma}+c_{L}\right)h_{n}}(u_{n}^{x})^{\left(c_{\gamma}+c_{L}\right)h_{n}}\mathrm{d}s$
$\leq\frac{\text{cst}}{\eta}\sup_{u\in[-1,1]}\left(\frac{g(x-h_{n}u)}{g(x)}\right)(u_{n}^{x})^{\left(c_{\gamma}+c_{L}\right)h_{n}}\int_{t}^{\infty}s^{-1/\gamma(x)+\varepsilon-1+\left(c_{\gamma}+c_{L}\right)h_{n}}\mathrm{d}s,$
for $\varepsilon\in(0,1/\gamma(x))$, using Potter’s bound. Let $\varepsilon^{\prime}\in(0,1/\gamma(x)-\varepsilon)$ be given. It is possible to find $n^{\prime}$ such that $\left(c_{\gamma}+c_{L}\right)h_{n}\leq\varepsilon^{\prime}$ for all $n\geq n^{\prime}$. Hence for $n\geq n^{\prime}$ we have that the integrand $s\mapsto s^{-1/\gamma(x)+\varepsilon-1+\left(c_{\gamma}+c_{L}\right)h_{n}}$ is integrable and consequently
@@
$\leq\limsup_{n\to\infty}\left\{\frac{\text{cst}}{\eta}\sup_{u\in[-1,1]}\left(\frac{g(x-h_{n}u)}{g(x)}\right)(u_{n}^{x})^{\left(c_{\gamma}+c_{L}\right)h_{n}}\int_{t}^{\infty}s^{-1/\gamma(x)+\varepsilon-1+\left(c_{\gamma}+c_{L}\right)h_{n}}\mathrm{d}s\right\}$
$=\mathcal{O}\left(t^{-1/\gamma(x)+\varepsilon+\varepsilon^{\prime}}\right),$
@@ A.4 Proof of Theorem 4 (Functional CLT)
are mutually independent and each have the same distribution as the first original sample block $\{(X_{1},Y_{1}),\ldots,(X_{r_{n}},Y_{r_{n}})\}$.
@@
$\widetilde{V}^{{\dagger},x}_{n}(s)=\frac{1}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\sum_{i=1}^{m_{n}}\sum_{j=1}^{r_{n}}K\left(\frac{x-X^{{\dagger}}_{(i-1)r_{n}+j}}{h_{n}}\right)1_{\left\{Y^{{\dagger}}_{(i-1)r_{n}+j}>su_{n}^{x}\right\}},$
@@
$\lim_{n\to\infty}\operatorname{cov}{\left(\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(s),\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(t)\right)}=\left(s\vee t\right)^{-1/\gamma(x)}g(x)\int K^{2}(u)\mathrm{d}u.$ (13)
@@
$\operatorname{cov}{\left(\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(s),\widetilde{\mathbb{V}}^{{\dagger},x}_{n}(t)\right)}=\frac{m_{n}r_{n}}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\biggl[\mathbb{E}\left[1_{\left\{Y_{0}>\left(s\vee t\right)u_{n}^{x}\right\}}K^{2}\left(\frac{x-X_{0}}{h_{n}}\right)\right]$
$\quad\quad-r_{n}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}\right]\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}^{x}\right\}}\right]$
$\quad\quad+\sum_{j=1}^{r_{n}}\left(1-\frac{j}{r_{n}}\right)\biggl\{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\right]$
$\qquad\qquad\qquad\qquad\qquad+\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]\biggr\}\biggr].$
@@
$\lim_{n\to\infty}\frac{m_{n}r_{n}}{n}\frac{\mathbb{E}\left[1_{\left\{Y_{0}>\left(s\vee t\right)u_{n}^{x}\right\}}K^{2}\left(\frac{x-X_{0}}{h_{n}}\right)\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}=\left(s\vee t\right)^{-1/\gamma(x)}g(x)\int K^{2}\left(u\right)\mathrm{d}u,$
which is the asymptotic covariance. It remains to show that the second and third terms disappear asymptotically. The second term vanishes by the assumption $\lim_{n\to\infty}{r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})}=0$. Since $m_{n}r_{n}/n=\mathcal{O}(1)$, the leading term in the third term is asymptotically proportional to
$\frac{1}{h_{n}\overline{F}^{x}(u_{n}^{x})}\sum_{j=1}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right].$
@@
$\limsup_{n\to\infty}\frac{1}{h_{n}\overline{F}^{x}(u_{n}^{x})}\sum_{j=1}^{m-1}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]$
$=\limsup_{n\to\infty}\sum_{j=1}^{m-1}h_{n}\int_{[-1,1]^{2}}\biggl\{\frac{\mathbb{E}\left[1_{\left\{Y_{0}>tu_{n}^{x}\right\}}1_{\left\{Y_{j}>su_{n}^{x}\right\}}\mid X_{0}=x+w_{1}h_{n},X_{j}=x+w_{2}h_{n}\right]}{\overline{F}^{x}(u_{n}^{x})}$
$\qquad\qquad\qquad\qquad\qquad\qquad\times g(x+w_{1}h_{n},x+w_{2}h_{n})K\left(w_{1}\right)K\left(w_{2}\right)\biggr\}\mathrm{d}(w_{1},w_{2})=0.$
@@
$\limsup_{n\to\infty}\frac{1}{h_{n}\overline{F}^{x}(u_{n}^{x})}\sum_{j=m}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>tu_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]\leq\varepsilon.$
@@
Assume conditions 1-41-2. Then the finite dimensional distributions of $\widetilde{\mathbb{V}}^{{\dagger},x}_{n}$ converge to those of a Gaussian process $\mathbb{V}^{x}$ with covariance function defined in (13).
@@
$S_{n,i}^{x}(s)=\sum_{j=1}^{r_{n}}K\left(\frac{x-X_{(i-1)r_{n}+j}}{h_{n}}\right)1_{\left\{Y_{(i-1)r_{n}+j}>su_{n}^{x}\right\}},\quad\overline{S}_{n,i}^{x}(s)=S_{n,i}^{x}(s)-\mathbb{E}\left[S_{n,1}^{x}(s)\right],$
$S_{n,i}^{{\dagger},x}(s)=\sum_{j=1}^{r_{n}}K\left(\frac{x-X^{{\dagger}}_{(i-1)r_{n}+j}}{h_{n}}\right)1_{\left\{Y^{{\dagger}}_{(i-1)r_{n}+j}>su_{n}^{x}\right\}},\quad\overline{S}_{n,i}^{{\dagger},x}(s)=S_{n,i}^{{\dagger},x}(s)-\mathbb{E}\left[S_{n,1}^{x}(s)\right],$
@@
where the summands are i.i.d. In the calculations below we show that the conditions of the Lindeberg–Lévy CLT are fulfilled. By Lemma 6 it only remains to show the negligibility condition. Since $\overline{S}_{n,1}^{x}(s)$ and $\overline{S}_{n,1}^{{\dagger},x}(s)$ have the same distribution, we may establish it for the former. Note that $m_{n}/(nh_{n}\overline{F}^{x}(u_{n}^{x}))=\mathcal{O}((r_{n}h_{n}\overline{F}^{x}(u_{n}^{x}))^{-1})$. Thus the negligibility condition thus follows if the expectation below is $o(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x}))$:
@@
$\mathbb{E}\left[\left(\sum_{j=1}^{r_{n}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right)^{2}1_{\left\{\lvert\overline{S}_{n,1}^{x}(s)\rvert>\varepsilon\sqrt{k_{n}h_{n}}\right\}}\right]$
$\quad+\left(r_{n}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}\right]\right)^{2}\mathrm{Pr}{\left(\left\lvert\overline{S}_{n,1}^{x}(s)\right\rvert>\varepsilon\sqrt{k_{n}h_{n}}\right)}$
$\quad-2r_{n}\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}\right]\sum_{j=1}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}1_{\left\{\lvert\overline{S}_{n,1}^{x}(s)\rvert>\varepsilon\sqrt{k_{n}h_{n}}\right\}}\right].$
@@
$r_{n}^{2}\ \mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}\right]^{2}=r_{n}^{2}\mathcal{O}\left(\left(h_{n}\overline{F}^{x}(u_{n}^{x})\right)^{2}\right)=o\left(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})\right).$
@@
$I_{1}$ $\leq\frac{1}{\varepsilon\sqrt{k_{n}h_{n}}}\sum_{j=1}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)^{2}1_{\left\{Y_{j}>su_{n}^{x}\right\}}\left\lvert\overline{S}_{n,1}^{x}(s)\right\rvert\right$]
$\leq\frac{1}{\varepsilon\sqrt{k_{n}h_{n}}}\sum_{j=1}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)^{2}1_{\left\{Y_{j}>su_{n}^{x}\right\}}\left(S_{n,1}^{x}(s)+\mathbb{E}\left[S_{n,1}^{x}(s)\right]\right)\right]$
@@
$\frac{\mathcal{O}\left(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})\right)}{\varepsilon\sqrt{k_{n}h_{n}}}\sum_{j=1}^{r_{n}}\mathbb{E}\left[K\left(\frac{x-X_{j}}{h_{n}}\right)^{2}1_{\left\{Y_{j}>su_{n}^{x}\right\}}\right]=\frac{\mathcal{O}\left(\left(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})\right)^{2}\right)}{\varepsilon\sqrt{k_{n}h_{n}}}=o\left(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})\right),$